# FedCache 2.0: Exploiting the Potential of Distilled Data in Knowledge Cache-driven Federated Learning

### v1

@@ 5.2 Performance Evaluation
Average User Model Accuracy. Table 1 displays the comparison of average UA on image recognition datasets, CIFAR10, CIFAR100, and CINIC10, with two degrees of data heterogeneity, . As displayed, FedCache 2.0 significantly outperforms all considered state-of-the-art methods across both model homogeneous and heterogeneous settings, demonstrating its superior performance and robustness in diverse edge scenarios. This substantial improvement is attributed to the enriched information characterization provided by distilled data and effective personalized optimization facilitated by device-centric cache sampling. Additional evaluations on audio understanding and mobile sensor data mining are provided in Appendix G.
@@
Communication Cost. Figure 3 illustrates the learning curves for image recognition tasks, plotting average UA against communication cost. As shown, FedCache 2.0 exhibits significantly steeper convergence curves, reaching acceptable average UA more efficiently than competing methods, regardless of data heterogeneity, model structures, and datasets. This indicates that FedCache 2.0 can achieve robust performance improvement with reduced communication overhead, making it suitable for deployment in resource-constrained edge environments with limited wireless bandwidth. The reduction in communication cost is attributed to FedCache 2.0’s elimination of transferring cumbersome model parameters between devices and the server. Alternatively, FedCache 2.0 leverages compact distilled data as knowledge to facilitate communication-efficient personalized optimization on devices. Quantitative communication information and experimental results on additional application tasks are provided in Appendix F and Appendix G, respectively.
@@
5.3 Ablation Study
Impact of Cache Sampling Strategy. Table 3 presents the average UA with different values. We can conclude that increasing in the early stage generally improves the average UA due to the richer information provided by a greater number of cached samples. However, the performance gain diminishes as approaches 1. This decline is likely due to the introduction of data distribution bias across devices, which is harmful to the system’s performance.
Impact of Model Settings. Table 3 presents the average UA for different model configurations across image recognition datasets. The results indicate that heterogeneous model settings yield higher average UA compared to homogeneous settings constrained by the weakest end devices. The improvements stem from the support of more powerful devices to deploy larger and more complex models, which can make full use of computational resources among heterogeneous devices to achieve better performance. These findings underscore the benefits of model heterogeneity flexibility in FedCache 2.0.
@@ 6 Discussion
Broader Impacts. FedCache 2.0 introduces tolerance for uncertain connection, allowing devices to engage in FEL at their convenience, without relying on any two devices being online at the same time. This feature is particularly advantageous in dynamic network environments, especially for IoT applications prone to unstable connectivity from power outages or limited signal coverage. Moreover, FedCache 2.0’s generalized data anonymization broadens its utility across various data modalities and application tasks. For instance, FedCache can be seamlessly integrated into smart healthcare and e-commerce recommendation systems, facilitating personalized trained models’ deployment on smartwatches or mobile phones to monitor users’ health status and shopping preferences.
@@
Limitations. In terms of potential limitations of FedCache 2.0, devices may maliciously upload misleading or poisoned distilled data to the server, which could negatively affect the overall system performance. In addition, the dataset distillation process conducted on devices demands considerable computational resources. This request can somewhat lead to slower training procedures on devices with low hardware capabilities or those constrained by battery life.
@@ 7 Conclusion
Section D provides information on the model configurations for both homogeneous and heterogeneous settings, data partition strategies, and hyper-parameter settings across the experiments.
Section E introduces the methodology for quantifying the communication costs of FL algorithms, including specific measures for different types of transmitted information like model weights, logits, sample index, and distilled data.
Section F elaborates on detailed quantitative communication costs and efficiency speed-up ratios for image recognition tasks with different degrees of data heterogeneity.
Section G provides additional evaluations of FedCache 2.0 on audio understanding and mobile sensor data mining tasks. It presents the average user model accuracy and communication efficiency for these tasks, demonstrating the effectiveness of FedCache 2.0 in diverse data modalities.
@@ Appendix A Privacy Demonstration of Federated Dataset Distillation
In this paper, federated data distillation is employed to develop abstract semantic representations on individual devices collaboratively. Figure 4 illustrates the comparison of raw local data and distilled data on three image recognition datasets. As shown in Figure 4, raw datasets on each client feature rich and identifiable visual information, posing severe risks to user privacy. For instance, in the CIFAR-10 cifar10 dataset, the unprocessed images vividly display the characteristic features of certain common objects such as cars, birds, and plains. Similarly, in the CINIC-10 cinic10 dataset, the distinct contours of people are clearly discernible. On the contrary, the images become significantly obscured and unidentifiable after distillation. Specifically, the distilled data across all datasets exhibit a substantial reduction in distinct object contours and detailed features, preserving only rudimentary aspects of color and shape. While the transformed data preserves enough statistical attributes for effective local model training, it lacks the granularity necessary for image recovery. Hence, this transformation greatly enhances user privacy by significantly mitigating the risk of personal data leakage when such distilled data is processed centrally on servers.
@@ Appendix B Illustration of Tasks and Datasets
Image Recognition. Image recognition is a fundamental task in computer vision, involving the assignment of images into predefined classes based on their visual content, such as color, texture, and shape. In our experiments, we evaluate our method on three image recognition datasets, which are CIFAR-10, CIFAR-100 cifar10, and CINIC-10 cinic10. All of these datasets consist of common objects in daily life, represented by vehicles, animals, and humans. They offer a diverse and challenging set of images that test the robustness and precision of FL algorithms in handling visual data.
@@
Audio Understanding. Audio understanding involves classifying audio clips into various categories based on their acoustic properties. We employ the UrbanSound8K dataset urbansound in our experiments, which contains labeled sound excerpts from 10 distinct classes, including car horns, street music, and children playing. Each sound clip is less than four seconds long, providing a compact yet comprehensive acoustic profile for analysis. This dataset is particularly valuable for evaluating the performance of audio recognition FL systems in naturalistic urban environments.
Mobile Sensor Data Mining. Mobile sensor data mining focuses on analyzing data from mobile device sensors to infer user activities or contextual settings. In our experiments, we utilize the TMD dataset carpineti2018custom designed to identify transportation modes based on smartphone sensor readings. The TMD dataset is ideal for assessing the capability of FL algorithms to process and interpret complex, real-time sensor data in the context of mobile computing.
@@ Appendix C Detailed Information on Model Structures
In this section, we provide detailed descriptions of the model structures used in our experiments. For experiments on image recognition, we employ deep residual network (ResNet) he2016deep with different numbers of layers to represent models with different structures, which are ResNet-T, ResNet-S, ResNet-M, and ResNet-T. For experiments on audio understanding as well as mobile sensor data mining, we adopt fully connected networks (FCN) that satisfy task-specific inputs and outputs sizes. We provide detailed information on model structures in Figure 5, and the number of parameters of adopted models in Table 4.
Appendix D Implementation Details
In this section, we outline the implementation details of our experiments, elaborating on the model configurations for both homogeneous and heterogeneous settings, the adopted data partition strategy, and the hyper-parameters employed across our experiments.
Model Configurations. We conduct experiments in both model-homogeneous and model-heterogeneous settings, respectively. For image recognition experiments with homogeneous models, all clients employ ResNet-L. For image recognition experiments with heterogeneous models, models on clients are evenly distributed among ResNet-S, ResNet-M, and ResNet-L. Note that for this set of experiments, we only consider FedKD wu2022communication and FedCache wu2024fedcache as baseline algorithms to compare with. This choice is based on the fact that MTFL mills2021multi, kNN-Per marfoq2022personalized, and SCDPFL chen2024spectral support only homogeneous models across clients. In addition, ResNet-T is adopted to facilitate the interaction of model parameters between clients and the server within the FedKD algorithm in all image recognition experiments. For audio understanding on UrbanSound8K urbansound and mobile sensor data mining on TMD carpineti2018custom, we exclusively employ the model-homogeneous setting, with all clients deploying a uniform FCN model, whose structures are denoted as FCN-U and FCN-T, respectively.
Method | Source of Code | Hyper-Parameters | Value |
MTFL | mtflurl | optimizer | customized |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
server_lr | 0.01 | ||
bn_private | usyb | ||
0.9 | |||
0.999 | |||
0.5 | |||
kNN-Per | knnperurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
aggregator_type | centralized | ||
client_type | KNNPerClient | ||
SCDPFL | scdpflurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 2 | ||
communicaion round | 100 | ||
0.4 | |||
0.3 | |||
FedKD | fedkdurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
tmax | 0.98 | ||
tmin | 0.95 | ||
FedCache | fedcacheurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
1.5 | |||
16 | |||
FedCache 2.0 | Our Implementation | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 5 | ||
communicaion round | 15 | ||
distillation learning rate | 0.001 | ||
0.5 |
Data Partition. For all experiments conducted in this paper, we utilize the data partition strategy provided by FedML he2020fedml to simulate decentralized datasets for clients. We adopt the hyper-parameter to control the degree of data heterogeneity. A smaller number of results in a greater degree of heterogeneity among clients. Notably, the training and testing datasets for the same client have identical distributions, whereas different clients typically have different data distributions, as illustrated in Figure 6.
Hyper-parameter Settings. We provide a default setting in Table 5, which is implemented in all experiments unless stated otherwise. Note that the table only showcases a subset of hyperparameters for baseline algorithms; those not mentioned retain their original settings as specified in the corresponding open-source code. Exclusively, we uniformly set the learning rate to 0.1 and the communication round of kNN-Per to 200 for experiments on mobile sensor data mining.
Method | CIFAR10/45% | CINIC10/40% | CIFAR100/30% | |
MTFL | N/A | N/A | N/A | |
kNN-Per | N/A | N/A | N/A | |
SCDPFL | 17.4G (1.0) | 7.8G (1.0) | 17.6G (1.0) | |
FedKD | N/A | N/A | N/A | |
FedCache | N/A | N/A | N/A | |
FedCache 2.0 | 389.1M (45.8) | 193.8M (41.2) | 609.3M (29.6) | |
Method | CIFAR10/30% | CINIC10/30% | CIFAR100/15% | |
MTFL | N/A | 39.9G (1.0) | 64.5G (1.0) | |
kNN-Per | 12.1G (1.6) | N/A | 17.6G (3.7) | |
SCDPFL | 19.4G (1.0) | 10.6G (3.8) | 15.9G (4.1) | |
FedKD | N/A | N/A | N/A | |
FedCache | N/A | N/A | N/A | |
FedCache 2.0 | 315.9M (62.9) | 473.5M (86.3) | 197.5M (334.4) |
Appendix E Communication Cost Calculation
@@
MTFL. Both model and optimizer parameters are serialized and transmitted in each communication round. All parameters are encoded as tensors in the float format, with each element occupying 4 bytes.
kNN-Per and SCDPFL. Similar to MTFL, these methods involve transmitting model parameters in float tensors in each communication round.
FedKD. Parameters of the student model (ResNet-T) are transmitted in each round.
FedCache. This method introduces additional complexity by transmitting sample hashes (in float format), sample index (in integer format), and logits (in float format).
@@
Appendix F Quantitative Communication Information for Image Recognition
In this section, we provide detailed quantitative communication information for all considered FL algorithms in image recognition tasks with homogeneous models. We measure the communication cost in terms of the amount of information transmitted between clients and the server until the average User model Accuracy (UA) reaches a specified threshold. We also provide the communication efficiency speed-up ratio, which compares the communication cost of each method to that of the least efficient baseline method achieving the same threshold precision. Table F quantitates the corresponding efficiency of FedCache 2.0 compared with baseline algorithms. It is evident that FedCache 2.0 significantly reduces the communication cost compared to other state-of-the-art methods, with 28.6 to 332.4 communication efficiency improvements compared with state-of-the-art methods.
Method | ||
MTFL | 56.5 | 46.9 |
kNN-Per | 54.3 | 54.1 |
FedCache 2.0 | 69.4 | 64.8 |
Method | ||
MTFL | 54.8 | 46.5 |
kNN-Per | 61.6 | 48.6 |
FedCache 2.0 | 73.3 | 61.7 |
Method | /50% | /45% |
MTFL | 13.0G (1.0) | 18.4G (1.0) |
kNN-Per | 7.5G (1.73) | 7.2G (2.6) |
FedCache 2.0 | 14.3M (930.9) | 19.1M (986.5) |
Method | /50% | /45% |
MTFL | 20.5G (1.0) | 11.9G (1.0) |
kNN-Per | 12.4G (1.7) | 11.0G (1.1) |
FedCache 2.0 | 14.8M (1418.4) | 15.4M (791.3) |
Appendix G Results on Additional Application Tasks
In this section, we evaluate the effectiveness of FedCache 2.0 across additional application tasks, specifically focusing on audio understanding and mobile sensor data mining. These tasks are critical for various real-world edge applications and offer broader evaluations of FedCache 2.0 across diverse data modalities.
G.1 Evaluation on Audio Understanding
Table 8 presents the average UA on UrbanSound8K dataset under two degrees of data heterogeneity. As observed, FedCache 2.0 consistently outperforms the baseline methods, achieving higher UA across both settings. Additional experimental results on communication efficiency are provided in Table 10 and Figure 8, demonstrating significant reductions in communication cost while maintaining superior performance.
G.2 Evaluation on Mobile Sensor Data Mining
Table 8 presents the average UA on TMD dataset under two degrees of data heterogeneity. Results demonstrate that FedCache 2.0 significantly outperforms existing methods under both degrees of data heterogeneity. We also provide qualitative and quantitative evaluation of FedCache 2.0 in Figure 8 and Table 10, respectively, indicating that FedCache 2.0 not only achieves higher system performance but also requires significantly less communication overhead compared to baseline methods.
[heading=subbibliography]

### v2

@@ 4.3 Evaluation on Image Classification
Average User Model Accuracy. Table V displays the comparison of average UA on image recognition datasets, CIFAR10, CIFAR100, and CINIC10, with two degrees of data heterogeneity, . As displayed, FedCache 2.0 significantly outperforms all considered state-of-the-art methods across both model homogeneous and heterogeneous settings, demonstrating its superior performance and robustness in diverse edge scenarios. This substantial improvement is attributed to the enriched information characterization provided by distilled data and effective personalized optimization facilitated by device-centric cache sampling.
| Method | ||
| MTFL | 56.5 | 46.9 |
| kNN-Per | 54.3 | 54.1 |
| FedCache 2.0 | 69.4 | 64.8 |
| Method | /50% | /45% |
| MTFL | 13.0G (1.0) | 18.4G (1.0) |
| kNN-Per | 7.5G (1.73) | 7.2G (2.6) |
| FedCache 2.0 | 14.3M (930.9) | 19.1M (986.5) |
@@
Communication Cost. Figure 5 illustrates the learning curves for image recognition tasks, plotting average UA against communication cost. As shown, FedCache 2.0 exhibits significantly steeper convergence curves, reaching acceptable average UA more efficiently than competing methods, regardless of data heterogeneity, model structures, and datasets. This indicates that FedCache 2.0 can achieve robust performance improvement with reduced communication overhead, making it suitable for deployment in resource-constrained edge environments with limited wireless bandwidth. The reduction in communication cost is attributed to FedCache 2.0’s elimination of transferring cumbersome model parameters between devices and the server. Alternatively, FedCache 2.0 leverages compact distilled data as knowledge to facilitate communication-efficient personalized optimization on devices. We also provide detailed quantitative communication burden for all considered FL algorithms in image recognition tasks with homogeneous models. We measure the communication cost in terms of the amount of information transmitted between clients and the server until the average User model Accuracy (UA) reaches a specified threshold. We also provide the communication efficiency speed-up ratio, which compares the communication cost of each method to that of the least efficient baseline method achieving the same threshold precision. Table V quantitates the communication cost and efficiency speed-up ratio of FedCache 2.0 compared with baseline algorithms. It is evident that FedCache 2.0 significantly reduces the communication cost to reach a given target average UA compared to baseline algorithms, with 28.6 to 332.4 times of communication efficiency improvements compared with state-of-the-art methods.
| Method | ||
| MTFL | 54.8 | 46.5 |
| kNN-Per | 61.6 | 48.6 |
| FedCache 2.0 | 73.3 | 61.7 |
| Method | /50% | /45% |
| MTFL | 20.5G (1.0) | 11.9G (1.0) |
| kNN-Per | 12.4G (1.7) | 11.0G (1.1) |
| FedCache 2.0 | 14.8M (1418.4) | 15.4M (791.3) |
![[Uncaptioned image]](x7.png)
4.4 Evaluation on Audio Understanding
Table VIII presents the average UA on UrbanSound8K [42] dataset under two degrees of data heterogeneity. As observed, FedCache 2.0 consistently outperforms the baseline methods, achieving higher UA across both settings. Additional experimental results on communication efficiency are provided in Table VIII and Figure VIII, demonstrating significant reductions in communication cost while maintaining superior performance.
4.5 Evaluation on Mobile Sensor Data Mining
Table XI presents the average UA on TMD [39] dataset under two degrees of data heterogeneity. Results demonstrate that FedCache 2.0 significantly outperforms considered baseline algorithms regardless of data heterogeneity settings. We also provide qualitative and quantitative evaluation of FedCache 2.0 in Figure XI and Table XI, respectively, indicating that FedCache 2.0 not only achieves higher system performance but also requires significantly less communication overhead.
4.6 Ablation Study
Impact of Cache Sampling Strategy. Table XII presents the average UA with different values. We can conclude that increasing in the early stage generally improves the average UA due to the richer information provided by a greater number of cached samples. However, the performance gain diminishes as approaches 1. This decline is likely due to the introduction of data distribution bias across devices, which is harmful to the system’s performance.
Impact of Model Settings. Table XIII presents the average UA for different model configurations across image recognition datasets. The results indicate that heterogeneous model settings yield higher average UA compared to homogeneous settings constrained by the weakest end devices. The improvements stem from the support of more powerful devices to deploy larger and more complex models, which can make full use of computational resources among heterogeneous devices to achieve better performance. These findings underscore the benefits of model heterogeneity flexibility in FedCache 2.0.
@@ 5.1 Broader Impacts
FedCache 2.0 introduces tolerance for uncertain connection, allowing devices to engage in FEL at their convenience, without relying on any two devices being online at the same time. This feature is particularly advantageous in dynamic network environments, especially for IoT applications prone to unstable connectivity from power outages or limited signal coverage. Moreover, FedCache 2.0’s generalized data anonymization broadens its utility across various data modalities and application tasks. For instance, FedCache can be seamlessly integrated into smart healthcare and e-commerce recommendation systems, facilitating personalized trained models’ deployment on smartwatches or mobile phones to monitor users’ health status and shopping preferences.
5.2 Limitations
@@
One of the primary concerns with FedCache 2.0 is the vulnerability to malicious participants. Since devices upload distilled data to the server, there is a risk that malicious devices could inject poisoned distilled data into the system. This could distort the knowledge cache and degrade the performance of the entire system. Another limitation of FedCache 2.0 is the computational burden imposed by the dataset distillation process on devices. Federated dataset distillation introduces additional optimization burdens besides local model training, demanding considerable computational resources. This request can somewhat lead to slower training procedures on devices with low hardware capabilities or those constrained by battery life.
6 Related Work
6.1 Federated Learning with Dual Learning and Model Personalization
FL methods are designed with the motivation to adapt to various hardware configurations through model personalization, while also addressing diverse user preferences via learning personalization. A variety of approaches have been developed to tackle the dual personalization challenges within FL. Differentiated client-side model optimization objectives are implemented in studies such as [17, 45, 11, 22, 29, 46], enabling trained models to generalize across clients with varying local data distributions. Novel client-server interaction designs, which depart from the traditional FedAvg [26], are explored in [16, 47, 19, 21, 48] to better accommodate diverse client hardware configurations with differently structured models. Furthermore, hybrid approaches such as [25, 12] are proposed to simultaneously address both model and learning personalization in FL.
Different from prior works, FedCache 2.0 implements a customized knowledge cache that stores and organizes distilled data to assist local model training, allowing heterogeneous clients to benefit from personalized knowledge fetched from the remote knowledge cache with efficient communication.
6.2 Federated Learning in Edge Computing
The efficiency of executing FL at the network edge has become a hot topic. Research works such as [49, 50, 51, 52] investigate the technical frameworks required for running FL algorithms on devices constrained by computational power or storage resources. Device heterogeneity and connection uncertainty in edge environments are tackled by methodologies such as [14, 25, 53, 12]. Moreover, [54, 55, 56, 57, 58] extends FL to a multi-tier architecture involving end-edge-cloud collaborations, enhancing model training efficiency and final performance by leveraging the edge as a bridge between devices and the cloud during the training process.
FedCache 2.0 is well-suited for edge environments as it effectively addresses key challenges such as limited bandwidth, device heterogeneity, and uncertain connectivity. By caching distilled data on the server and conducting device-centric cache sampling, FedCache 2.0 achieves state-of-the-art performance with reduced communication overhead, while ensuring that even resource-constrained devices can efficiently utilize global knowledge for locally adapted model training.
6.3 Federated Learning with Alternative Information
Instead of transmitting model parameters, alternative information is utilized in the FL training process by a series of recent works. Model-agnostic outputs are exchanged between clients and the server in [25, 59, 48, 21, 20], allowing deployment of customized on-device models across resource-heterogeneous clients. Additionally, methodologies involving uploading mixed or distilled data from clients to servers are proposed in [60, 61], significantly reducing communication overhead while maintaining client data privacy.
FedCache 2.0 extends these ideas by leveraging distilled data as the medium of knowledge exchange. Through implementing knowledge caching with federated dataset distillation, FedCache 2.0 provides a more performance-guaranteed paradigm for information transfer between clients and the server, while ensuring efficient communication.
@@ 7 Conclusion
Section D introduces the methodology for quantifying the communication costs of FL algorithms, including specific measures for different types of transmitted information like model weights, logits, sample index, and distilled data.
@@ Appendix A Privacy Demonstration of Federated Dataset Distillation
In this paper, federated data distillation is employed to develop abstract semantic representations on individual devices collaboratively. Figure 6 illustrates the comparison of raw local data and distilled data on three image recognition datasets. As shown in Figure 6, raw datasets on each client feature rich and identifiable visual information, posing severe risks to user privacy. For instance, in the CIFAR-10 [40] dataset, the unprocessed images vividly display the characteristic features of certain common objects such as cars, birds, and plains. Similarly, in the CINIC-10 [41] dataset, the distinct contours of people are clearly discernible. On the contrary, the images become significantly obscured and unidentifiable after distillation. Specifically, the distilled data across all datasets exhibit a substantial reduction in distinct object contours and detailed features, preserving only rudimentary aspects of color and shape. While the transformed data preserves enough statistical attributes for effective local model training, it lacks the granularity necessary for image recovery. Hence, this transformation greatly enhances user privacy by significantly mitigating the risk of personal data leakage when such distilled data is processed centrally on servers.
@@ Appendix B Illustration of Tasks and Datasets
Image Recognition. Image recognition is a fundamental task in computer vision, involving the assignment of images into predefined classes based on their visual content, such as color, texture, and shape. In our experiments, we evaluate our method on three image recognition datasets, which are CIFAR-10, CIFAR-100 [40], and CINIC-10 [41]. All of these datasets consist of common objects in daily life, represented by vehicles, animals, and humans. They offer a diverse and challenging set of images that test the robustness and precision of FL algorithms in handling visual data.
@@
Audio Understanding. Audio understanding involves classifying audio clips into various categories based on their acoustic properties. We employ the UrbanSound8K dataset [42] in our experiments, which contains labeled sound excerpts from 10 distinct classes, including car horns, street music, and children playing. Each sound clip is less than four seconds long, providing a compact yet comprehensive acoustic profile for analysis. This dataset is particularly valuable for evaluating the performance of audio recognition FL systems in naturalistic urban environments.
Mobile Sensor Data Mining. Mobile sensor data mining focuses on analyzing data from mobile device sensors to infer user activities or contextual settings. In our experiments, we utilize the TMD dataset [39] designed to identify transportation modes based on smartphone sensor readings. The TMD dataset is ideal for assessing the capability of FL algorithms to process and interpret complex, real-time sensor data in the context of mobile computing.
@@ Appendix C Detailed Information on Model Structures
In this section, we provide detailed descriptions of the model structures used in our experiments. For experiments on image recognition, we employ deep residual network (ResNet) [44] with different numbers of layers to represent models with different structures, which are ResNet-T, ResNet-S, ResNet-M, and ResNet-T. For experiments on audio understanding as well as mobile sensor data mining, we adopt fully connected networks (FCN) that satisfy task-specific inputs and outputs sizes. We provide detailed information on model structures in Figure 7, and the number of parameters of adopted models in Table XIV.
Appendix D Communication Cost Calculation
@@
MTFL [11]. Both model and optimizer parameters are serialized and transmitted in each communication round. All parameters are encoded as tensors in the float format, with each element occupying 4 bytes.
kNN-Per [17] and SCDPFL [18]. Similar to MTFL, these methods involve transmitting model parameters in float tensors in each communication round.
FedKD [16]. Parameters of the student model (ResNet-T) are transmitted in each round.
FedCache [25]. This method introduces additional complexity by transmitting sample hashes (in float format), sample index (in integer format), and logits (in float format).
@@
![[Uncaptioned image]](extracted/5924587/pqy.jpg)
![[Uncaptioned image]](extracted/5924587/ss.jpg)
![[Uncaptioned image]](extracted/5924587/wuzhiyuan.jpg)
![[Uncaptioned image]](extracted/5924587/wyw.jpg)
![[Uncaptioned image]](extracted/5924587/lm.jpg)
![[Uncaptioned image]](extracted/5924587/gb.png)
![[Uncaptioned image]](extracted/5924587/wjy.jpg)